OpenStudy (anonymous):

What is the solution to the equation? 6x – 27 = –3 A. x = –5 B. x = –4 C. x = 4 D. x = 5

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

D

OpenStudy (anonymous):

@TheAsker2002

OpenStudy (anonymous):

Can you explain?

OpenStudy (anonymous):

You Can use mathway for an explanation

OpenStudy (kidrah69):

Substitute the X values in place of X in the equation

OpenStudy (anonymous):

I used a paper i cant type the entire thing i did

OpenStudy (anonymous):

ok

OpenStudy (kidrah69):

@bleebleebleed please avoid giving direct answers people are here for HELP not ANSWERS

OpenStudy (anonymous):

Please do not type to me in that tone of voice

OpenStudy (kidrah69):

Ok sweetie :) i did that so it would stand out

OpenStudy (anonymous):

Well, It did Kinda Lol

OpenStudy (anonymous):

@kidrah69 So now i know how to make my reply stand out :P

OpenStudy (kidrah69):

Please stop spamming the post.

OpenStudy (anonymous):

@kidrah69 he isnt

TheSmartOne (thesmartone):

Back to the topic, add 27 on both sides. What do you get?

OpenStudy (anonymous):

Which sides?

TheSmartOne (thesmartone):

6x – 27 = –3 +27 +27 -------------

OpenStudy (anonymous):

54 and 24

OpenStudy (anonymous):

Oh let me tag @Sam_Aka_Sara

TheSmartOne (thesmartone):

Remeber there is a negative in front of the 27. You could rewrite what I wrote above as: \(\sf 6x-27+27 = -3+27\)

OpenStudy (anonymous):

I thought it was a minus

TheSmartOne (thesmartone):

-27 + 27 = 27 - 27 = ?

OpenStudy (anonymous):

0

TheSmartOne (thesmartone):

and -3 + 27 = 27 -3 = ?

OpenStudy (anonymous):

27 - 3?

OpenStudy (anonymous):

24

TheSmartOne (thesmartone):

So now you get: \(\sf 6x = 24\) What is 24/6 = ?

TheSmartOne (thesmartone):

also btw, \(\sf -3+27= +27 -3\) if we re-arrange the terms we can get that and it's much easier subtraction :)

TheSmartOne (thesmartone):

Whatever you get as 24/6 = ? is your answer :)

OpenStudy (anonymous):

Thank you so much!

TheSmartOne (thesmartone):

Anytime :D